Find the measure of an angle whose supplement measures 6 degrees more than 4 times its complement.
Answer by mananth(16949)

let angle be x
complement = 90-x
supplement =180-x
180-x=4(90-x)+6
180-x=360-4x+6
3x=360-180+6
3x=186
x=186/3
x=62 deg
...
m.ananth@hotmail.ca